Find the EAR in each of the following cases Stated Rate (APR) Number of Times Compounded Effective Rate (EAR) 11.50 % Quarterly % 14.50 % Monthly % 17.50 % Daily % 13.50 % Semiannually %
Answer #1

EAR = (1+r)^n - 1 Compounded Quarterly EAR= EAR= (1 +11.50%/4)^4 - 1 12.01% Compounded Monthly EAR= EAR= (1 +14.50%/12)^12 -

EAR Quarterly: 12.01% ;

Monthly: 15.50%

Daily : 19.12%

Semi-Annually: 13.96%